# Computational electro-optic frequency comb spectroscopy

### v1

@@ 3 Experimental results
Fig 3b 2 3 4 31 0.3 0.019
Fig 3d 1 3 4 0.4 41 0.4 0.016
Fig 5 1 5.1 1.50 25 4.8 0.020
Fig 6 0.5 5 1.52 0.2 51 10 0.050
Fig 7 0.5
5 (PM1)
4.9 (PM2)
1.64 (PM1)
1.68 (PM2)
51 10 0.015
4 Discussion and conclusions
In this paper, we have presented a computational method for spectroscopic measurements based on a rapidly reconfigurable EO comb generator fed by a narrow-linewidth cw laser. The capability of this system to deal with different classes of spectroscopic samples has been demonstrated by numerical simulations. Despite limitations imposed by the available electronic equipment, the experimental results obtained in our laboratory confirm the feasibility of the proposed approach. We have successfully reconstructed the transmission of an optical bandpass filter with a FWHM below the resolution of a typical grating-based optical spectrum analyzer (<0.05<0.05 nm). We have also reconstructed a molecular absorption line of HCN with a frequency sampling of $500$ MHz covering >10>10 GHz. For this particular measurement, we have used two phase modulators in cascade (to increase the number of attainable comb lines). However, this two-stage configuration could be avoided with an ultra-low-$V_{\pi}$ phase modulator capable of handling high RF power [38]. The number of system reconfigurations has been chosen to keep the acquisition time within $10$ ms, although shorter times can be considered as long as the sampling spectra span over a sufficiently broad optical bandwidth (see Section S4 of the Supplementary Material). Additionally, the measurement speed could be further increased with faster reconfigurable electronics.
Compared with traditional dual-comb or interferometric techniques, our method reduces spectroscopic system complexity while preserving the high spectral sampling of frequency combs. In our experiments, we employed a highly stable ultra-narrow-linewidth laser (<0.1<0.1 kHz). Nevertheless, given the line spacings considered (hundreds of MHz to a few GHz), more economical sources with broader linewidths, such as DFB lasers, could still yield comb spectra reasonably close to the ideal noise-free case. Although optical signal amplification was not required in this work, in practical scenarios with significant losses it may become necessary. In such cases, the perturbation introduced by amplified spontaneous emission noise in the sampling spectra would need to be evaluated, and the truncated inverse technique might need to be replaced or complemented by more advanced reconstruction algorithms [43].
@@ 4 Discussion and conclusions
We have focused on phase modulation and showed that a wide range of comb spectra can be produced using multitone RF signals. The potential extension to more general EO comb configurations—including both phase and intensity modulators—is theoretically straightforward. Such multi-stage generators could increase several times the number of comb lines (and, hence, the bandwidth) beyond the maximum value demonstrated here ($M\leq 27$ within 20 dB of power). Additionally, a Mach-Zehnder modulator (MZM) driven by an AWG could be used to map custom-designed digitally generated spectra into the optical domain [31]. These extensions to other configurations—or even a change in the comb platform—are possible provided that two basic requirements are met: the ability to dynamically reconfigure the comb generator in a predictable or measurable manner, and the generation of sufficiently dissimilar and stable combs. Experimentally, the sampling spectra may be measured directly—for a fixed driving configuration—using heterodyne interferometry [46] or a high-resolution spectrum analyzer.
@@ Acknowledgments
Appendix A. Multi-tone driving signal
@@
which corresponds to equation (3), where we have omitted the subscript $i$ to simplify the notation. Accordingly, the field $E$ can be expressed as:
@@
where $E_{0}$ is the field amplitude, $\nu_{0}$ is the carrier frequency, $j$ denotes the imaginary unit, $f_{s}$ is the fundamental RF frequency and $\beta_{l}$ and $\varphi_{l}$ are, respectively, the modulation index and the phase of the l-harmonic. This equation corresponds to Eq. (2) in Sec. (II).
@@
$\tilde{E}(\nu)=$ $E_{0}\sum_{k_{1}=-\infty}^{\infty}\sum_{k_{2}=-\infty}^{\infty}\dotsb\sum_{k_{h}=-\infty}^{\infty}j^{k_{1}+k_{2}+\dotsb+k_{h}}$ (14)
@@
$\cdot e^{\,j\left(k_{1}\varphi_{1}+k_{2}\varphi_{2}+\dotsb+k_{h}\varphi_{h}\right)}$
@@
$J_{n}(x,y;e^{j\phi})=\sum_{l=-\infty}^{\infty}\left(e^{j\phi}\right)^{l}J_{n-2l}(x)J_{l}(y).\\$ (16)
@@
The relationship between GBFs of order $n$ and n-n is given by [34]:
@@
Appendix B. Experimental setup for dual-comb spectroscopy
The experimental setup used to perform dual-comb spectroscopy is shown in Fig. 8. The output power ($20$ mV) from a narrow-linewidth laser (Koheras Basik E15, NKT) is split into two arms, each containing an identical phase modulator (MPZ-LN-10, IXblue) driven by an amplified RF signal coming from a frequency generator. The two driving RF signals have slightly different frequencies $(f_{s}=1$ GHz and $f_{s}+\delta f_{s}=1.0001$ GHz). An acousto-optic frequency shifter (AA Opto-electronic), placed in one arm, shifts the laser frequency by $f_{AO}=80$ MHz to prevent overlap in the down-converted RF spectrum. The two generated optical combs are recombined in fiber and then split again into two paths. In one of them, the light passes through a spectroscopic sample before detection, while the other path contains only a photodetector and serves as a reference. This scheme—known as symmetric or collinear, as both combs pass through the sample—measures only the sample absorption. Although the comb phase information is not directly accessible in this configuration (it requires additional processing and an extended RF bandwidth [53]), the optical setup remains robust to external perturbations. Balanced photodetectors are used at the end of both paths, BPD1 (PDB470C, Thorlabs) and BPD2 (PD100B, Koheron), and the detection signals are digitalized by a single oscilloscope (MSO8104, Rigol). We acquire $20$ ms traces, each containing $2000$ consecutive interferograms. The Fourier transform of each trace yields an RF comb with a line spacing of $100$ kHz. The powers applied to the phase modulators are adjusted to minimize amplitude variations in the dual-comb RF spectrum, at the cost of reducing the total number of resolved comb lines to a maximum of $13$. The sample response, encoded in the RF comb measured in the sample arm, is retrieved by normalizing to the spectrum obtained in the reference arm. With the sole exception of the sample itself, all components in the dual-comb setup use polarization-maintaining fiber. Three RF generators are used in the setup: two for driving the phase modulators (DSG821, Rigol; SSG3032X, Siglent), and one for the AOFS (SDG6032X, Siglent). All generators, along with the oscilloscope, are locked to a highly stable reference clock (OCXO B08, Rigol).
@@ Appendix B. Experimental setup for dual-comb spectroscopy
Appendix C. Inverse problem for the molecular absorption line
Let $x$ denote our spectroscopic sample, which is the discretized version of a molecular transition line as shown in Fig. 9(a) (typically described by a Voigt profile). Over the considered optical bandwidth, we will assume that the maximum value of the transmission $T(\nu)$ as a function of the optical frequency $\nu$ is $1$. In this case, due to the characteristics of our probe comb, we reformulate the inverse problem to reconstruct $x^{\prime}=1-T$, instead of $T$, see Fig. 9(b). After discretization, we obtain $\mathbf{Ax^{\prime}=b^{\prime}}$, where $\mathbf{A}$ is the square matrix $M\times M$ formed by the consecutive probe combs arranged in rows and $\mathbf{b^{\prime}}$ is the vector of measured power values. According to the definition of $x^{\prime}$}, we can write $\mathbf{b^{\prime}}$ as $\mathbf{b^{\prime}=A(1_{M}-x)}$, where $1_{M}$ is a column vector consisting of $M$ ones, and $\mathbf{x}$ is the transmission $T$ at each comb line. Assuming that the total power of all comb lines sums to $1$, and using $\mathbf{Ax=b}$ (see Eq.(10)), the i-th component of $b^{\prime}$ is simply given by $b^{\prime}_{i}=1-b_{i}$. If, instead, the plateau transmission is $b_{0}$, the normalized value of $b^{\prime}_{i}$ is given by $b^{\prime}_{i}=1-b_{i}/b_{0}$. The value of $b_{0}$ can be determined in several ways, as discussed in the main text and in Ref. [41].
Figure 9: (a) Transmission curve $T(\nu)$ for a molecular transition. It typically follows a Voigt profile, reaching a minimum value $T_{min}$. (b) In our computational approach, we reformulate the inverse problem to retrieve $1-T(\nu)$.

### v2

@@ III EXPERIMENTAL RESULTS
Fig. 3(b) 2 3 4 31 0.3 0.018
Fig. 3(d) 1 3 4 0.4 41 0.4 0.016
Fig. 5 1 5.1 1.50 25 4.8 0.018
Fig. 6 0.5 5 1.52 0.2 51 10 0.070
Fig. 7 0.5 5 (PM1) 4.9 (PM2) 1.64 (PM1) 1.68 (PM2) 51 10 0.009
IV DISCUSSION AND CONCLUSIONS
In this paper, we have presented a computational method for spectroscopic measurements based on a rapidly reconfigurable EO comb generator fed by a narrow-linewidth cw laser. The capability of this system to deal with different classes of spectroscopic samples has been demonstrated by numerical simulations. Despite limitations imposed by the available electronic equipment, the experimental results obtained in our laboratory confirm the feasibility of the proposed approach. We have successfully reconstructed the transmission of an optical bandpass filter with a FWHM below the resolution of a typical grating-based optical spectrum analyzer (<0.05<0.05 nm). We have also reconstructed a molecular absorption line of HCN with a frequency sampling of $500$ MHz covering >10>10 GHz. For this particular measurement, we have used two phase modulators in cascade (to increase the number of attainable comb lines). However, this two-stage configuration could be avoided with an ultra-low-$V_{\pi}$ phase modulator capable of handling high RF power [38].The number of system reconfigurations was chosen to keep the acquisition time within 10 ms, although this number could be reduced to speed up acquisition, as long as the generated spectra cover a sufficiently broad optical bandwidth and the system to be solved remains square (see Section S4 of the Supplementary Material). Additionally, the measurement speed could be further increased with faster reconfigurable electronics.
Compared with traditional dual-comb or interferometric techniques, our method reduces spectroscopic system complexity while preserving the high spectral sampling of frequency combs. In our experiments, we employed a highly stable ultra-narrow-linewidth laser (<0.1<0.1 kHz). Nevertheless, given the line spacings considered (hundreds of MHz to a few GHz), more economical sources with broader linewidths, such as distributed feedback (DFB) lasers, could still yield comb spectra reasonably close to the ideal noise-free case. Although optical signal amplification was not required in this work, in practical scenarios with significant losses it may become necessary. In such cases, the perturbation introduced by amplified spontaneous emission noise in the sampling spectra would need to be evaluated, and the truncated inverse technique might need to be replaced or complemented by more advanced reconstruction algorithms [43].
@@ IV DISCUSSION AND CONCLUSIONS
We have focused on phase modulation and showed that a wide range of comb spectra can be produced using multitone RF signals. The potential extension to more general EO comb configurations—including both phase and intensity modulators—is theoretically straightforward. Such multi-stage generators could increase several times the number of comb lines (and, hence, the bandwidth) beyond the maximum value demonstrated here ($M\leq 27$ within 20 dB of power). Additionally, a Mach-Zehnder modulator (MZM) driven by an AWG could be used to map custom-designed digitally generated spectra into the optical domain [31]. These extensions to other configurations—or even a change in the comb platform—are possible provided that two basic requirements are met: the ability to dynamically reconfigure the comb generator in a predictable or measurable manner, and the generation of sufficiently dissimilar and stable combs. Experimentally, the sampling spectra may be measured directly—for a fixed driving configuration—using heterodyne interferometry [46] or a ultra-high-resolution spectrum analyzer.
@@ Acknowledgments
Appendix A Multi-tone driving signal
@@
which corresponds to equation (3), where $f_{s}$ is the fundamental RF frequency and $\beta_{l}$ and $\varphi_{l}$ are, respectively, the modulation index and the phase of the l-harmonic. We have omitted the subscript $i$ to simplify the notation. Accordingly, the field $E$ can be expressed as:
@@
where $E_{0}$ is the field amplitude, $\nu_{0}$ is the carrier frequency and $j$ denotes the imaginary unit. This equation corresponds to Eq. (2) in Sec. (II).
@@
$\tilde{E}(\nu)=$ $E_{0}\sum_{k_{1}=-\infty}^{\infty}\sum_{k_{2}=-\infty}^{\infty}\dotsb\sum_{k_{h}=-\infty}^{\infty}e^{\,j\left(k_{1}\varphi_{1}+k_{2}\varphi_{2}+\dotsb+k_{h}\varphi_{h}\right)}$ (14)
@@
$J_{k}(x,y;e^{j\phi})=\sum_{l=-\infty}^{\infty}\left(e^{j\phi}\right)^{l}J_{k-2l}(x)J_{l}(y).\\$ (16)
@@
The relationship between GBFs of order $k$ and k-k is given by [34]:
@@
Appendix B Experimental setup for dual-comb spectroscopy
The experimental setup used for dual-comb spectroscopy is shown in Fig. 8. The output power ($20$ mV) from a narrow-linewidth laser (Koheras Basik E15, NKT) is split into two arms, each containing an identical phase modulator (MPZ-LN-10, iXblue) driven by an amplified RF signal coming from a frequency generator. The two driving RF signals have slightly different frequencies $(f_{s}=1$ GHz and $f_{s}+\delta f_{s}=1.0001$ GHz). An acousto-optic frequency shifter (AA Opto-Electronic), placed in one arm, shifts the laser frequency by $f_{AO}=80$ MHz to prevent overlap in the down-converted RF spectrum. The two generated optical combs are recombined in fiber and then split again into two paths. In one of them, the light passes through a spectroscopic sample before detection, while the other path contains only a photodetector and serves as a reference. This scheme—known as symmetric or collinear, as both combs pass through the sample—enables the measurement of the sample absorption. Although phase information is not directly accessible in this configuration (it requires additional processing and an extended RF bandwidth [53]), the optical setup remains robust to external perturbations. Balanced photodetectors are used at the end of both paths, BPD1 (PDB470C, Thorlabs) and BPD2 (PD100B, Koheron), and the detection signals are digitalized by a single oscilloscope (MSO8104, Rigol). We acquire $20$ ms traces, each containing $2000$ consecutive interferograms. The Fourier transform of each trace yields an RF comb with a line spacing of $100$ kHz. The powers applied to the phase modulators are adjusted to minimize amplitude variations in the dual-comb RF spectrum, at the cost of reducing the total number of resolved comb lines to a maximum of $13$. The sample response, encoded in the RF comb measured in the sample arm, is retrieved by normalizing to the spectrum obtained in the reference arm. With the sole exception of the sample itself, all components in the dual-comb setup use polarization-maintaining fiber. Three RF generators are used in the setup: two for driving the phase modulators (DSG821, Rigol; SSG3032X, Siglent), and one for the AOFS (SDG6032X, Siglent). All generators, along with the oscilloscope, are locked to a highly stable reference clock (OCXO B08, Rigol).
@@ Appendix B Experimental setup for dual-comb spectroscopy
Appendix C Inverse problem for the molecular absorption line
Let $x$ denote our spectroscopic sample, which is the discretized version of a molecular transition line as shown in Fig. 9(a) (typically described by a Voigt profile). Over the considered optical bandwidth, we will assume that the maximum value of the transmission $T(\nu)$ as a function of the optical frequency $\nu$ is $1$. In this case, due to the characteristics of our probe comb, we reformulate the inverse problem to reconstruct $x^{\prime}=1-T$, instead of $T$, see Fig. 9(b). After discretization, we obtain $\mathbf{Ax^{\prime}=b^{\prime}}$, where $\mathbf{A}$ is the square matrix $M\times M$ formed by the consecutive probe combs arranged in rows and the vectors $\mathbf{x^{\prime}}$ and $\mathbf{b^{\prime}}$ are, respectively, the sample response and the set of measured power values. According to the definition of $x^{\prime}$}, we can write $\mathbf{b^{\prime}}$ as $\mathbf{b^{\prime}=A(1_{M}-x)}$, where $1_{M}$ is a column vector consisting of $M$ ones, and $\mathbf{x}$ is the transmission $T$ at each comb line. Assuming that the total power of all comb lines sums to $1$, and using $\mathbf{Ax=b}$ (see Eq.(10)), the i-th component of $\mathbf{b^{\prime}}$ is simply given by $b^{\prime}_{i}=1-b_{i}$. If, instead, the plateau transmission is $b_{0}$, the normalized value of $b^{\prime}_{i}$ is given by $b^{\prime}_{i}=1-b_{i}/b_{0}$. The constant $b_{0}$ can be determined in several ways, as discussed in the main text and in Ref. [41].
Figure 9: Inverse problem for reconstructing the absorption line. (a) Transmission curve $T(\nu)$ for a molecular transition. It typically follows a Voigt profile, reaching a minimum value $T_{min}$. (b) In our computational approach, we reformulate the inverse problem to retrieve $1-T(\nu)$.